OpenStudy (anonymous):

How do I figure out 5 times the square root of 320 and simplify I'm in seventh grade so I need clear details

OpenStudy (jdoe0001):

 \(\bf 5 \sqrt{320} \implies 5 \sqrt{2\times2\times2\times2\times2\times2\times5}\\ 5\sqrt{(2^3)^2 \times 5} \implies 5 \sqrt{2^2\times 2^2 \times 2^2 \times 5} \implies 5 \times 2 \times2 \times2 \sqrt{5}\\ 40\sqrt{5}\)